### v1

# Invisible Labor in Open Source Software Ecosystems

Abstract.
Invisible labor is work that is not fully visible, not appropriately compensated, or both. In open source software (OSS) ecosystems, essential tasks that do not involve code (like content moderation) often become invisible to the detriment of individuals and organizations. However, invisible labor is so difficult to measure that we do not know how much of OSS activities are invisible. Our study addresses this challenge, demonstrating that roughly half of OSS work is invisible. We do this by developing a survey technique with cognitive anchoring that measures OSS developer self-assessments of labor visibility and attribution. Survey respondents ($n=142$) reported that their work is more likely to be nonvisible or partially visible (i.e. visible to at most 1 other person) than fully visible (i.e. visible to 2 or more people). Furthermore, cognitively anchoring participants to the idea of high work visibility increased perceptions of labor visibility and decreased visibility importance compared to anchoring to low work visibility. This suggests that advertising OSS activities as “open” may not make labor visible to most people, but rather lead contributors to overestimate labor visibility. We therefore add to a growing body of evidence that designing systems that recognize all kinds of labor as legitimate contributions is likely to improve fairness in software development while providing greater transparency into work designs that help organizations and communities achieve their goals.
labor, feminism, openness, survey, articulation work, infrastructuring, contribution, anchoring
ccs: Software and its engineering Open source modelccs: Software and its engineering Maintaining softwareccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Computer supported cooperative workccs: Social and professional topics Employment issuesccs: Social and professional topics Computing organizationsccs: Social and professional topics Project management techniquesccs: Social and professional topics Software maintenance
1. Introduction
Figure 1. (a) Invisible labor is labor that is not fully visible and/or undercompensated. Labor is not fully visible when labor data doesn’t exist, isn’t shared, or isn’t accessible. Labor is undercompensated when it doesn’t receive enough credit, pay, or new opportunities, among others forms of compensation. (b) Our study demonstrates that roughly two thirds of labor in OSS is not visible, either by not being visible or only shared with one other person. We also found that about half of OSS labor activities do not receive credit, as demonstrated here by the 9 in 10 individuals who do not receive credit from some fraction of projects they contribute to.
(a) Invisible labor becomes so when it is not visible (because it doesn’t exist, isn’t shared, or isn’t accessible) and/or when labor is undercompensated (because it doesn’t receive credit, doesn’t receive pay, doesn’t facilitate new opportunities, etc.). (b) One third of labor in our study was visible, one third partially visible, and one third invisible, and more than half of labor did not receive credit.
Figure 1. (a) Invisible labor is labor that is not fully visible and/or undercompensated. Labor is not fully visible when labor data doesn’t exist, isn’t shared, or isn’t accessible. Labor is undercompensated when it doesn’t receive enough credit, pay, or new opportunities, among others forms of compensation. (b) Our study demonstrates that roughly two thirds of labor in OSS is not visible, either by not being visible or only shared with one other person. We also found that about half of OSS labor activities do not receive credit, as demonstrated here by the 9 in 10 individuals who do not receive credit from some fraction of projects they contribute to.
The processes of creating and maintaining open source software (OSS) involve many diverse tasks beyond just contributing code (Giuri et al., 2010; Eghbal, 2016; Casari et al., 2021; Trinkenreich et al., 2020; Geiger et al., 2021). Conference organizing, community management, finance management, interaction design, data entry, and documenting have all become essential to OSS production and repair (Gousios et al., 2008; Eghbal, 2016, 2020; D’Ignazio and Klein, 2020; Ramin et al., 2020; Young et al., 2021; Trinkenreich et al., 2020). Despite growing interest in supporting noncode contributions (e.g. National Science Foundation, 2023), much of this labor remains undocumented, unattributed to individuals, and “practically invisible” to outsiders and newcomers (Young et al., 2021; Geiger et al., 2021). This often makes noncode contributions forms of invisible labor, undercompensated or uncompensated activities that are “rendered invisible…because they take place out of sight, or because they lack physical form altogether” (D’Ignazio and Klein, 2020).
Labor invisibility creates significant challenges for individuals and organizations involved in OSS. For individuals, lower task visibility makes it difficult to demonstrate the value of their contributions (Suchman, 1995; D’Ignazio and Klein, 2020). In turn, these individuals receive less compensation, whether in terms of future opportunities (Riehle, 2007; Babcock et al., 2017), pay (Toxtli et al., 2021), or social status (Stewart, 2005). And so these individuals frequently burn out and stop performing crucial activities (Eghbal, 2020) that a growing number of scientists, organizations, and governments depend on (Eghbal, 2016). Furthermore, these individuals are often members of minoritized groups (Frluckaj et al., 2022), thereby reinforcing demographic representation disparities (Babcock et al., 2017; D’Ignazio and Klein, 2020).
For organizations and governments, labor invisibility limits work transparency and therefore task measurement, coordination, and control (Suchman, 1995). This limits organizations’ knowledge into whom they fund in open source, what those individuals do, and how those activities shape valued outcomes (Casari et al., 2021). And although invisibility has some benefits for workers in these organizing contexts (like anonymity), computing researchers have long argued that invisibility prevents equity more than creates it (Suchman, 1995; Star and Strauss, 1999; D’Ignazio and Klein, 2020). Indeed, the reliance of open source ecosystems on volunteerism necessitates creating communities where volunteers find benefit in performing all kinds of activities (i.e. not just coding) (Fang and Neufeld, 2009; von Krogh et al., 2012; Eghbal, 2020).
How to make OSS labor (more) visible then? We likely need to (a) recognize diverse types of work that individuals perform and (b) measure how much of those activities people perform. On the former, task taxonomies have seen modest success. For example, academics have started adopting contributorship taxonomies like CRediT to increase attribution and transparency (Brand et al., 2015; Larivière et al., 2016; Holcombe, 2019). Likewise, software developers are experimenting with manual and automated attribution systems (e.g. All Contributors and octohatrack, respectively) with some success (Young et al., 2021). Measuring amounts of labor, though, has proven more elusive, particularly in OSS. Many automated measurement techniques (like GitHub’s contribution activity visualizations) only measure code as “contributions,” thereby conferring lower status to noncoding activities. This provides limited, often overly reductive assessments of “how much got done” (D’Ignazio and Klein, 2020; Young et al., 2021) and can be less meaningful to workers than manual attribution (Monroy-Hernández et al., 2011). And while computational solutions hold potential for spaces like crowd work (Toxtli et al., 2021), they seem unlikely to pass muster in open source ecosystems given the variety of computational and noncomputational interaction contexts involved (coding, conferences, direct chats, etc.).
To address this gap, this work measures invisible labor in OSS ecosystems through two contributions. First, we introduce a technique for measuring labor visibility in open source software development. We use a participatory feminist approach (Bardzell, 2010; D’Ignazio and Klein, 2020) — namely, asking workers directly — to understand their expert perceptions of both how often their work is visible to others and how visible it is to others. We do this by asking developers how often their work is visible to 2 or more people (“fully visible”), to 1 other person (“partially visible”), and to no other people (“nonvisible”). Controlling whether participants see these questions starting from high visibility ($\geq 2$, 1, 0) or low visibility (0, 1, $\geq 2$) with survey randomization (Tourangeau et al., 2013) enables us to estimate labor visibility via cognitive anchoring (Kahneman et al., 1982; Tourangeau, 2018).
Second, we estimate invisible labor in OSS ecosystems by measuring labor attribution alongside our visibility measurement technique. We anonymously surveyed 142 open source software developers across varied fields — from scientific computing to app development — about credit they receive for OSS tasks, their satisfaction with that credit, and the visibility of their work. Randomly giving half of the respondents the descending or ascending orders measured the visibility of the labor they perform and how cognitive anchoring shapes their perceptions. Our results show that across fields roughly half of OSS efforts do not receive credit. OSS activities are more likely nonvisible or partially visible (visible to at most 1 other person) than fully visible (to 2 or more people). Furthermore, individuals anchored to full visibility tended to view their own work as more visible and feel more satisfied with the credit they receive than individuals primed with nonvisibility. This suggests that people in ecosystems with norms of full visibility may overestimate the visibility of their work, making attribution, compensation, and governance inherently more difficult.
We begin by defining labor visibility, invisible labor, and their roles in Computer-Supported Cooperative Work (CSCW). We then describe our methodology and results from a global sample of software developers. We close by discussing what our findings mean for individuals and organizations involved in OSS along with design considerations for overcoming these challenges. Fig. 1 overviews our findings.
2. Background
To measure something that, by implication, one “cannot see,” it helps to understand what it means for work to be “visible” and “invisible” in the context of OSS. In short, labor becomes invisible when (1) labor data is not fully visible because it does not exist, is not shared, or is not accessible, and (2) labor is undercompensated. The following sections address these concepts in turn.
2.1. Labor Visibility
When we call work “visible,” we typically do not mean that we can physically “see” the work as it happens. Rather, visibility is a metaphor that captures the availability of information about work, approval of that information for sharing, and accessibility of that information (Stohl et al., 2016; Poster et al., 2016). In labor contexts, availability means that information about the tasks being performed has been created and stored somewhere (e.g. via recordkeeping). Then, those records need to be shared, whether compelled by the law, social norms, or personal conviction, with someone other than the creator. Finally, labor information becomes accessible when an individual knows both where to find the shared information and has sufficient skill to understand what tasks have been performed. Information about labor must meet all three of these conditions — it must exist, be shared, and be accessible — in order for it to be fully visible (Stohl et al., 2016).
More common perhaps is partial visibility, wherein tasks are less documented, less shared, less reachable, or less understandable. Alternatively, tasks may be documented but information only shared, reachable, or understandable to certain individuals (e.g. to experienced group members, but not newcomers) (Fang and Neufeld, 2009). Partial visibility is rarely due to lack of competence; to the contrary, specialized work is often the least visible because experts insulate others from needing to know the steps involved (Daniels, 1987; Suchman, 1995).
Outside of OSS, academic labor taxonomies like CRediT (Allen et al., 2014; Brand et al., 2015) exhaustively list individuals’ activities because, even if a reader witnessed every action that researchers took, readers would have limited visibility into what it took to conduct a study. Likewise, OSS labor records (e.g. the Github event stream, octohatrack, and All Contributors) attribute artifact activities, education, management, maintenance, security, and other roles (Young et al., 2021). Unfortunately, these systems rarely provide consistent, standardized recordkeeping (Young et al., 2021) or accessible sharing of those records (Monroy-Hernández et al., 2011). This creates a need for new measurement techniques as it leaves much labor invisible.
2.2. Invisible Labor
Invisible labor has become a popular term to describe marginalized forms of work both inside and outside the home, like childcare (Daniels, 1987), data entry (D’Ignazio and Klein, 2020), and crowd work (Toxtli et al., 2021). In one sense, labor becomes invisible if information about the work being done does not exist, is not shared, cannot be reached, or cannot be understood. Any of these conditions is sufficient to make information about labor, and therefore the doing of labor, invisible (Stohl et al., 2016).
Also implied by the concept is that certain tasks tend to be uncompensated or undercompensated (Acker, 1973; Daniels, 1987; Babcock et al., 2017; D’Ignazio and Klein, 2020; Toxtli et al., 2021). Individuals who perform less-visible tasks may struggle to find financial remuneration (D’Ignazio and Klein, 2020; Toxtli et al., 2021), opportunities for promotion (Riehle, 2007; Babcock et al., 2017), and to gain social status in their communities (Stewart, 2005). These activities often fall to women, non-binary people, and other marginalized individuals as a product of gendered socialization processes (Daniels, 1987; Berdahl et al., 2018). This makes invisible labor particularly pernicious to those who seek greater equity in computing (D’Ignazio and Klein, 2020; Casari et al., 2021) and makes compensation the fourth determinant of labor invisibility. [^1] While others define invisible labor in the context of paid employment (Poster et al., 2016; Toxtli et al., 2021), they also include activities that are unpaid but expected and, more generally, conceive of invisible labor as “activities that are tied to a job and its rewards” and from which an “employer reaps profits” (Poster et al., 2016, p.7-8). In OSS, formal employers sometimes do and sometimes do not exist, but many parties benefit from OSS regardless (e.g. companies, governments, scientists). For that reason, we argue volunteer activities in OSS ecosystems still merit inclusion, thereby broadening our conceptualization of “pay” into “compensation.” Together, labor is invisible whenever it is not fully visible, not appropriately compensated, or both (Fig. 1a).
CSCW has long sought to make labor more visible. Star & Strauss (Strauss, 1985; Star, 1991; Star and Strauss, 1999) coined articulation work [^2] Articulation work is “work that gets things back ‘on track’ in the face of the unexpected [and] is invisible to rationalized models of work,” a classic example of support acts (Star, 1991). in part to raise the visibility of efforts to support distributed work through mechanism design (e.g. new organizing forms, technologies, procedures) (Schmidt and Bannon, 1992; Schmidt and Simone, 1996; Star and Strauss, 1999). More recently, Cherry (Cherry, 2016), Ludwig & colleagues (Ludwig et al., 2018), Gray & Suri (Gray and Suri, 2019), and D’Ignazio & Klein (D’Ignazio and Klein, 2020) revealed other instances of invisibility in computing. Content moderation, proof reading, e-waste processing, and countless digitally-tethered activities struggle with invisibility. In effort to measure invisible labor, Toxtli and colleagues (Toxtli et al., 2021) recently developed a browser plugin to measure invisible labor in crowd work, finding that Amazon Mechanical Turkers spend nearly a third of their time on invisible tasks.
Like with crowd work, OSS development and maintenance activities largely remain invisible (Suchman, 1995; Eghbal, 2016, 2020; Casari et al., 2021), with two consequences. First, labor that deserves compensation is not likely to receive it without greater visibility (Casari et al., 2021). Second, we cannot understand how OSS ecosystems function without greater visibility into the activities of OSS (Suchman, 1995; Casari et al., 2021). Lest we assume this only applies to noncoding activities, consider a coding example too: Projects often combine substantially differentiated coding activities, like bug fixing and feature creation, under the label “code contributions” (Casari et al., 2021). Even activities historically valued as “contributions” are not fully visible for lack of fidelity, thereby making it difficult to assess the health of projects. This further motivates the need for new techniques to measure invisible labor among both code and noncode OSS activities.
3. Methods
Within OSS ecosystems, manual and automated tools have had mixed success at measuring diverse activities, code and noncode alike. As an alternative, we used a participatory feminist technique: working with developers to design a survey, and then directly asking a global sampling of developers about their OSS activities. The following sections describe how we designed and conducted the survey.
3.1. Survey Instrument Design
We codesigned a Qualtrics web survey (Tourangeau et al., 2013) with practicing developer advocates (among the coauthors). Given the importances of visibility and compensation for invisible labor, we dedicated a portion of the survey to each. We iteratively developed the questions and their responses among ourselves and with our Institutional Review Board (IRB), which approved the design before the start of the study.
Broadly, the survey focused participant attention on recent collaborative OSS projects and the breadth of tasks performed therein. This gave us access to participant perceptions of tasks beyond code contributions while still including coding activities. We asked a series of questions about the credit they receive as one form of compensation. Then, we asked several questions about how often other people know about the tasks they perform to measure labor visibility. See Appendix A for the full survey.
3.1.1. Guiding Participant Attention
A participant’s experience with a survey begins with recruitment materials. These materials, along with the consent form (Q1) and other questions throughout the survey, needed to carefully avoid the term “invisible labor” because topics related to supportive labor tend to be femininely coded — and even denigrated — in many global contexts (Berdahl et al., 2018; D’Ignazio and Klein, 2020). Using the term “invisible labor” in a survey could incentivize gendered signaling of masculinity and femininity or political signaling. Instead, we devised other terms to help participants recall the full breadth of tasks that people perform in their projects. We advertised the survey around “receiving credit” for their work as this is an important concept to many in OSS communities (Young et al., 2021) but which should be less likely to evoke either self-selection into the survey based on the survey focus or gendered social desirability bias (Krumpal, 2013).
The survey began by prompting the participant to focus on their experiences with open source projects that they worked on with other people during the last two years (Q2). We limited the survey to this time period to reduce experience underretrieval due to excessive cognitive load (Lenzner et al., 2010; Tourangeau, 2018). Focusing their attention on collaborative open source projects then gave us access to OSS projects that pose the greatest challenge to sustainability (Eghbal, 2020).
Later, we focused the participant’s attention on how “tasks vary in shape, size, and how visible they are to other people” (Q5). We asked them to “think through the different kinds of tasks you have done as part of these projects over the past 2 years.” We did this to provide rich retrieval cues about varied tasks, which tend to improve recall (Tourangeau, 2018). This broadened their perception out from “just code” to include as many tasks as possible within their work, thereby bringing their visible and invisible labor to the fore. For those who already considered noncoding activities to be contributions, this would not affect their judgment; however, for those who might consider code the only legitimate type of “contribution” (Eghbal, 2020; Casari et al., 2021), this design avoided the term contribution (via “task” instead) while improving information retrieval about diverse tasks (Tourangeau, 2018).
3.1.2. Compensation Questions
We asked participants about the artifacts and use cases of their OSS collaborations (Q3-4). We did this to measure survey distribution reach, participant diversity, and compensation variation across those settings. After discussion with our IRB, we chose not to include other demographic questions to allow us to distribute the survey globally while protecting varied, nation-specific participant rights through anonymity.
This followed with several questions about receiving credit for tasks, which we defined for them as “the project told someone that you did something for the project” (Q6-10). We asked about how many projects they received credit from (via 5-point Likert scale), how many tasks they received credit for (Likert scale), the medium through which they received credit (multiple select), and their satisfaction with the credit they received (Likert scale). These questions helped us gauge how much compensation (here, attribution) individuals receive, how they receive it, and to measure any relationship strength between attribution and perceptions of visibility.
If we want participant responses to approximate frequencies, it becomes important to control for perceptual biases that result from different perceptions of Likert scale terms. Experts suggest linearizing response options (e.g. explicitly numbering response options with 1 through 5 in addition to response text) (Tourangeau et al., 2000) to overcome this. So, we linearized response options for all Likert scales (For example: 1 - Extremely dissatisfied, 2 - Dissatisfied, 3 - Neither satisfied nor dissatisfied, 4 - Satisfied, 5 - Extremely satisfied, I’m not sure). We discuss the visibility questions next.
3.1.3. Labor Visibility Questions
Unlike in crowd work where all work might take place on one platform (Toxtli et al., 2021), OSS activities take place across an ecosystem (Casari et al., 2021) that makes measurement via tracking difficult. We therefore measured labor visibility by asking participants how often different numbers of people knew about tasks they performed. This approximated how visible different tasks are to OSS communities. Behavioral frequency impressions (e.g. how often someone performs invisible tasks) are prone to under- and overestimation (Conrad et al., 1998). While we cannot directly verify their perceptions, our approach treats their perceptions as expertise (Bardzell, 2010). Combining randomization with human cognitive biases allowed us to estimate likely bounds on how much labor individuals perform at different visibility levels.
To that end, we began the labor visibility questions with a prompt (“Think about all the tasks you performed again,” Q11) to aid recall of their activities. Then, we presented participants with three questions:
  1. Q12.
    How often did 2 or more people know that you performed those tasks?
  2. Q13.
    How often did 1 other person know that you performed those tasks? [^3] We chose to use the phrasing “1 other person” instead of several more and less specific alternatives. We did this because in open source contexts some participants may believe that everyone’s work should be visible and therefore accessible. “Only 1 other person” and “just 1 other person” may invoke social desirability effects as participants may perceive the words as implying that their work should be more visible than it is, thereby shaming them into underreporting. Likewise, the word “exactly” provides a rich cue that might encourage participants to second guess themselves. However, using merely “1 person” comes at the risk of ambiguity as to whether or not cases in which $\geq 2$ people know about a task are subsets of cases in which 1 person knows about a task. We reasoned that “1 other person” provided sufficient specificity without risking social desirability bias. Furthermore, we enabled the “back” button in the survey so participants could revise their answers if they inferred a different meaning after viewing all three questions ($\geq 2$, 1, 0).
  3. Q14.
    How often did nobody else know that you performed those tasks?
Information presented earlier in a survey tends to produce assimilation effects — that is, later judgments are moved in the direction of earlier judgments (Tourangeau, 2018), reminiscent of the anchoring and adjustment heuristic or priming effects (Kahneman et al., 1982; Molden, 2014). So if we only presented Q12, Q13, and Q14 in that order (called $\geq 2$, 1, 0 hereafter for clarity), participant responses would have been likely to bias results toward presumptions of high visibility (2 or more people, or “openness”). Therefore, we randomized whether participants saw these questions in order of what we call anchoring to high visibility ($\geq 2$, 1, 0) or anchoring to low visibility (0, 1, $\geq 2$).
Following these, we asked participants how important it is to receive credit for tasks (Q16) along with several free response questions about what did and did not work well for them about how people give credit.
3.2. Study Procedures
To test the survey instrument and measure invisible labor in OSS, we recruited 142 participants between January and June, 2022. We did this by posting IRB-approved recruitment materials on the social media platform formerly known as Twitter with the hashtag #opensource. Retweets confirm that the recruitment messages reached open source participants on at least 5 continents (North America, South America, Europe, Africa, Australia) and a number of OSS developer communities (e.g. Python, Django, pan-African OSS, open science).
3.3. Validity & Reproducibility
We estimate that, very roughly, between 10 million and 100 million individuals participate in open source projects (Trujillo et al., 2022). With our sample of 142 participants, a 20% proportion percentage (based on five or more Likert scale options per question), and a 90% confidence interval, this gives a 5.5% margin of error for our survey.
While our results are not a formal psychometric scale, we calculated reliability estimates on the numeric measures of our survey to assess their internal consistency using coefficient omega total ($\omega_{total}$), now considered the most reliable measure, and Cronbach’s alpha ($\alpha$), a more traditional measure, along with our analyses (McNeish, 2018; Jebb et al., 2021). Reliability estimates of our numeric data suggest high internal consistency ($\omega_{Total}=0.84$, $\alpha=0.78$). For intrinsic consistency, contribution roles, survey data, and code for this project are available at the following:
https://github.com/meluso/open-source-credit-survey
4. Results
We performed four analyses with the survey data. We began by examining how many participants work on different artifacts and use cases to confirm diverse labor participation in our study. Then, we examined how often participants receive compensation (specifically credit) and how visible their work is. Last, we considered the relationships between variables and calculated causal relationships where possible.
4.1. Sample Captures Diverse OSS Labor Activities
Figure 2. Distributions of (a) artifacts and (b) use cases that survey participants contribute to. Questions were multiple select, so percents shown represent the fraction of participants who expressed that they work on the artifacts and use cases shown.
Horizontal bar graphs displaying distributions of the artifacts and use cases that participants contribute to. At least 30% of participants expressed that they work on each of libraries, applications, packages, web-related tools, and frameworks. A significant majority (68%) of respondents work on software infrastructure, and all other use cases are below 30%.
Figure 2. Distributions of (a) artifacts and (b) use cases that survey participants contribute to. Questions were multiple select, so percents shown represent the fraction of participants who expressed that they work on the artifacts and use cases shown.
Study participants reported working on a wide variety of software artifacts and use cases (Fig. 2). Our survey questions on artifacts and use cases (Q3 & Q4, respectively) each allowed participants to select multiple answers, so percents shown represent the proportion of participants involved in the respective categories.
Each of five categories were selected by more than 30% of the participants: software libraries, applications, packages, web-related tools, and frameworks. At least 20% of our sample reported contributing to each of four others including operating systems, scripts, data science, and programming languages. Hardware (15%) and something else (13%) saw lower figures, but at least 18 people selected each. Many of the “something else” responses consisted of documentation, community building, and research tools.
Project use cases were more skewed among participants. Two thirds (68%) reported working on software infrastructure, though most selected multiple responses (mean $=2.52$ responses, st. dev. $=1.87$). Six other response options received at least 20% including business & productivity applications, scientific research, content management, communication technologies, education, and browsers. Entertainment, publishing, manufacturing, and other uses cases all received less than 15%. Here, “something else” included developer tools, data science, and project management among others.
While our sample of 142 responses clearly cannot represent all OSS activity, this sample diversity suggests that trends in our data may reveal broader OSS labor trends.
4.2. Half of OSS Labor May Not Receive Attribution
Next, we examine the data from our compensation questions which focused on receipt of credit for work. The data suggest that participants receive credit inconsistently for the work they do (Fig. 3a-b) but that they still feel moderately satisfied, regardless of how much credit they receive (Fig. 3c-d).
Figure 3. Distributions of responses to questions about how often people receive credit (a) from specific projects and (b) for specific tasks; and distributions for how satisfied individuals are with (c) how often they receive credit and (d) the mediums through which they receive credit. Dots and error bars represent mean values with one standard deviation on the linearized scales. Colors show how individuals responded to (a), tracking those same individuals across the other questions.
The figure shows four horizontal bar graphs. (a) and (b) show wide distributions centered around “Some of them - 3”; (c) a slightly skewed distribution with most responses on “Satisfied - 4” and “Neither satisfied nor dissatisfied - 3”; and (d) a skewed distribution with most responses on “Satisfied - 4”.
Figure 3. Distributions of responses to questions about how often people receive credit (a) from specific projects and (b) for specific tasks; and distributions for how satisfied individuals are with (c) how often they receive credit and (d) the mediums through which they receive credit. Dots and error bars represent mean values with one standard deviation on the linearized scales. Colors show how individuals responded to (a), tracking those same individuals across the other questions.
Participants reported that they receive credit from approximately half of the projects they participate in (mean $=3.07$ on a 1 to 5 scale) and for half of the tasks they perform (mean $=2.96$, also a 1 to 5 scale). Significant variation existed among the responses though (st. dev. $=1.17$ for projects, st. dev. $=1.02$ for tasks). Ten percent or less reported receiving credit from projects and tasks either always or never.
In terms of frequency (Fig. 3c) and medium (Fig. 3d), participants reported feeling slightly satisfied with the credit they receive. Satisfaction with credit frequency skewed slightly left, falling just above “Neither satisfied nor dissatisfied” (mean $=3.36$, st. dev. $=1.02$). Visual inspection confirms that only about 1 in 5 respondents (19.0%) felt dissatisfied with how often they receive credit, while about 1 in 3 felt either lukewarm (31.6%) or satisfied (33.8%). Then, satisfaction with the medium through which individuals received credit skewed significantly left, falling closer to “Satisfied” than “Neither satisfied nor dissatisfied” (mean $=3.58$, st. dev. $=0.88$). Here, almost 1 in 2 (47.2%) reported feeling “Satisfied” while 1 in 4 (23.9%) answered “Neither satisfied nor dissatisfied.” Most of this credit came in the form of automated metrics (63.1%), project documentation (61.5%), social media posts (50.8%), and project memberships (47.7%).
Overall, nearly 9 in 10 (89.1%) participants reported that some of their work is not fully visible (and therefore invisible) while only 1 in 10 (10.9%) reported full visibility (Fig. 1b).
4.3. Anchoring Shifts Perceptions of Labor Visibility
Figure 4. Responses to “How often did {2 or more people, 1 other person, nobody else} know that you performed those tasks?” as described by response (a) means and (b-d) distributions. Participants who saw the questions in ascending order (0,1,$\geq 2$) tended to report that their work is more likely to be invisible (seen by nobody else) and less likely to be partially invisible (seen by 1 other person) than those who saw questions in the descending order ($\geq 2$,1,0). In (a) arrows indicate question viewing order, and errorbars show 95% confidence interval of difference between means.
(a) shows a plot of points with errorbars and connecting lines indicate the question viewing direction. (b-d) show vertical bar graphs of the distributions of responses to each of the two viewing directions.
Figure 4. Responses to “How often did {2 or more people, 1 other person, nobody else} know that you performed those tasks?” as described by response (a) means and (b-d) distributions. Participants who saw the questions in ascending order (0,1,$\geq 2$) tended to report that their work is more likely to be invisible (seen by nobody else) and less likely to be partially invisible (seen by 1 other person) than those who saw questions in the descending order ($\geq 2$,1,0). In (a) arrows indicate question viewing order, and errorbars show 95% confidence interval of difference between means.
Turning to labor visibility, participants either viewed questions starting from an anchor to high visibility (seen by $\geq 2$, 1, then 0 people) or low visibility (0, 1, $\geq 2$ people). Averaging the results of the two groups and placing them in proportion to one another (Fig. 1b) approximates how often labor is invisible (2 in 3, 66.7%) compared to fully visible (1 in 3, 33.3%). But separately, Fig. 4 shows that the anchors produced statistically significant differences for the two groups, both in their mean responses on each question (Fig. 4a) and the distributions of their responses (Fig. 4b-d). Together, their answers suggest ranges in which the “true” likelihood of labor visibility may fall.
The group anchored to high visibility (shown in purple) viewed questions in descending order. They usually reported that $\geq 2$ people knew about their work “Sometimes” or “Often” over the past two years (Fig. 4b); then that 1 other person would “Often” or “Always” know about their work (Fig. 4c); and finally that “Rarely” or “Sometimes” would nobody else know about their work (Fig. 4d). Again, survey respondents tend to answer later survey questions by making adjustments from earlier questions (Tourangeau, 2018). This progression suggests the group’s relative perceptions of the labor visibilities; to them, their work was most likely to be partially visible (1 person), less likely to be fully visible ($\geq 2$ people), and less likely still to be not visible (0 people). Cumulatively, the cases that guarantee invisibility (when labor is either not visible or partially visible) is more likely than full visibility.
The group anchored to low visibility (turquoise) significantly contrasted the previous group. Viewing questions in ascending order, they were more likely to report their work was “Often,” “Sometimes,” or “Rarely” visible to nobody else (Fig. 4d); “Sometimes” or “Often” visible to 1 other person (Fig. 4c); and “Sometimes” or “Rarely” again for $\geq 2$ people (Fig. 4b). Here too, this group likely made adjustments from their anchor (nonvisibility). Consequently, this group’s averages suggest that they perceive all three cases — full, partial, and nonvisibility — as happening roughly as often as one another, though partial visibility may occur slightly more often than nonvisibility based on their distinct 95% confidence intervals (Fig. 4a). Here, too, the cumulative combination of partially visible and nonvisible labor makes invisibility more likely than fully visible labor.
The two groups roughly agree on how often their OSS labor is fully visible, revealing that about 2 in 3 tasks are invisible (Fig. 1b). Still, the differences between groups on nonvisibility and partial visibility are significant, statistically and in implication. First, the group anchored to high visibility perceived invisible labor as less common than the low visibility anchored group, suggesting that anchoring may influence perceptions of labor visibility. This appears again in the difference between the partial visibility means: both groups perceived partial visibility as more likely than their respective anchors, but the groups differed on the likelihood depending on the initial framing. The qualitative differences between the response distributions in Figs. 4b & c further support this.
4.4. Visibility Anchoring May Shift Beliefs About the Importance of Credit
Figure 5. Correlations between numerical survey variables, ordered by clustering. Colored boxes are statistically significant ($p<0.01$) while gray boxes are not. Being seen by nobody else (I) was correlated with low scores on several other variables. Conversely, being seen by $\geq 2$ people (D) was often correlated with high scores on other variables. Being seen by 1 other person yielded mixed results.
A heatmap grid of correlation values with variables listed across the vertical and horizontal axes. More positively correlated relationships are colored increasingly darker shades of red; more negatively correlated relationships are colored increasingly darker shades of blue; uncorrelated relationships are colored white; and statistically insignificant relationships are light gray.
Figure 5. Correlations between numerical survey variables, ordered by clustering. Colored boxes are statistically significant ($p<0.01$) while gray boxes are not. Being seen by nobody else (I) was correlated with low scores on several other variables. Conversely, being seen by $\geq 2$ people (D) was often correlated with high scores on other variables. Being seen by 1 other person yielded mixed results.
Finally, we examine relationships between the quantitative variables of our survey. Calculating Pearson correlation coefficients between these variables reveals several trends (Fig. 5). Unsurprisingly, the likelihoods of receiving credit for specific tasks and from specific projects are highly positively correlated ($r=0.87$), as were participants’ satisfactions with the credit they receive ($r=0.73$, compare Fig. 3c & d). Credit frequency and participant satisfaction tended to correlate with one another (A, B, E, & F) and with full visibility (D). Partial visibility (C) also correlates with the former but less so, while negatively correlating with anchoring order (H, $r=-0.30$) and being not visible (I, $r=-0.27$). Indeed, negative correlations proved common for nonvisibility (A-F) and only positively correlated with anchoring to low visibility (H, $r=0.30)$. Anchoring to low visibility also correlated with importance of receiving credit (G, $r=0.26$), which in turn correlated with both credit frequencies ($r=0.23$ $r=0.20$, respectively).
Again, invisible labor consists of both visibility and compensation. To understand the relationship between these elements in OSS (here, with credit as compensation), we calculated a weighted average of the visibility measures (fully, partially, & nonvisible) and separately averaged the frequencies of receiving credit (credit for tasks, credit from projects). Fig. 6a shows the relationship between average visibility and average credit frequency for each participant ($r=0.53$). While this is correlative and not causative, visibility and credit clearly trend together. Few individuals’ responses significantly deviated from this trend. Calculating the average of averages for each dimension provides further confirmation that, in general, OSS labor is only partially visible (51.9%) and roughly half receives credit (50.1%).
The trend between visibility and credit becomes more important in the context of one last observation. Recall that the importance of receiving credit (Fig. 5, G) correlated with anchoring order (H). The importance of receiving credit (Q16) appeared after the visibility questions (Q12-14) meaning that anchoring order had the potential to yield causal effects for subsequent questions. Fig. 6b shows that anchoring had a statistically significant effect on how important credit was to participants. On average, participants anchored to low visibility thought receiving credit was $30.5\%$ more important than participants anchored to high visibility ($95\%$ confidence interval of ±$\pm 19.5\%$, $p<0.002$, Adj. $R^{2}=0.061$ with robust standard errors). We discuss the implications of this and our other findings next.
Figure 6. (a) Scatterplot of average visibility versus average credit frequency. Points indicate individual responses. The regression line shows the correlation between variables. (b) Linear regression of credit importance against visibility anchoring. Points indicate individual responses and are grouped on the x-axis by presentation order. The regression line of best fit shows a statistically significant positive relationship between labor invisibility (via anchoring) and greater importance of credit (Adj. $R^{2}=0.061$).
Two scatterplots, one of average labor visibility versus average credit frequency with a positive trend. The second regresses how important receiving credit is to participants against question presentation order. Here, the x-axis is broken out by presentation order (decreasing on the left, increasing on the right) and the y-axis shows importance responses from 1 (not important) to 5 (extremely important). A regression line of best fit is overlaid showing a slight slope upward from left to right.
Figure 6. (a) Scatterplot of average visibility versus average credit frequency. Points indicate individual responses. The regression line shows the correlation between variables. (b) Linear regression of credit importance against visibility anchoring. Points indicate individual responses and are grouped on the x-axis by presentation order. The regression line of best fit shows a statistically significant positive relationship between labor invisibility (via anchoring) and greater importance of credit (Adj. $R^{2}=0.061$).
5. Discussion
5.1. Findings & Implications
Labor becomes invisible labor when it is undercompensated, not fully visible, or both. Through a labor-diverse sample of OSS ecosystems, we found that invisible labor seems to pervade OSS ecosystems (Fig. 1b). Participants reported receiving credit for roughly half of their work. While this finding cannot speak to other forms of compensation (e.g. pay, social status, opportunities), credit is one of the least financially and socially expensive forms of compensation suggesting that this compensation rate is potentially an upper bound on compensation frequency.
Participants reported feeling only slightly satisfied with how often they receive credit and the mediums through which they receive it. This slight satisfaction suggests a possible explanation for why many individuals do not seek more credit. Many who were less satisfied with how much credit they receive listed sources of credit similar to those who were satisfied. However, satisfied individuals cited several sources that dissatisfied individuals cited less often including blogs (47.9% for satisfied, 5.3% for dissatisfied), presentations (37.5% vs. 10.5%), and project documentation (64.6% vs. 42.1%). It is possible, then, that these mediums might increase the efficacy of sharing information or the accessibility of information about OSS labor.
We then found that visibility anchoring shapes perceptions of the visibility of a person’s work. Asking participants about highly visible work (seen by $\geq 2$ people) first in our series of questions increased participant frequency perceptions of partially visible work (seen by 1 other person) and decreased frequency perceptions of low visibility work (seen by nobody else) compared to those whom we asked about low visibility first.
Our finding that visibility anchoring shifts labor visibility perceptions has two implications for OSS ecosystems. OSS communities tend to presume that work is visible because “openness means visibility” so to speak. This is not necessarily the case because visibility requires information to exist, get shared, and be accessible. As we have seen, OSS labor is only partially visible on average and roughly half does not receive credit. So presuming that openness creates visibility may actually bias OSS participants into perceiving invisible labor as less common than it is, thereby decreasing compensation efforts. The finding that visibility anchoring reduces credit importance perceptions further supports this likelihood. In addition, participants generally viewed partial visibility (seen by 1 other person) as more common than full visibility or nonvisibility. This confirms the precariousness of OSS labor noted by others (Eghbal, 2020) — here in the form of OSS labor visibility — for two reasons. First, if the singular external viewer of a task loses visibility into that task — by promotion, leaving the community, or even just forgetting — that labor becomes unknown to others. Second, some labor is already seen by nobody other than the worker. Either way, these tasks are invisible and therefore difficult for others to understand, compensate, and manage. Assuming this work is valuable to an OSS ecosystem, it reinforces the importance of accessible recordkeeping.
The correlation and regression results also imply that making OSS labor more visible probably requires giving individuals credit more often and more consistently. Yet even those who always receive credit did not report their work being completely visible. So though not all work should be fully visible, we join others who take the feminist position that all work should receive compensation (D’Ignazio and Klein, 2020). Invisible work — completely unknown to others — cannot receive compensation from anyone other than the individual performing the labor. If individuals have sufficient means (e.g. wealth, social status), they may be able to compensate themselves for performing invisible tasks, but most people cannot. We see this already in OSS when privileged individuals receive credit for activities that surpass community-defined thresholds of social status or centrality while less privileged individuals do not receive such credit (D’Ignazio and Klein, 2020; Young et al., 2021).
To remedy all of this, OSS ecosystems likely need more robust qualitative and quantitative documentation so that labor data exists; more effective dissemination practices so labor data gets shared; more easily understood data communication so labor data becomes accessible; and more credit in valued forms so all labor receives compensation. Treating all work as contributions may be an effective way to start (Casari et al., 2021; Allen et al., 2014).
5.2. Limitations & Future Work
Our work has several limitations. First of all, survey instruments tend to prove more reliable in the form of validated scales. Questions remain about how feasible creating a scale would be for something so difficult to trace which motivated the exploratory nature of this study. That said, existing information visibility scales (ter Hoeven et al., 2021) may guide the labor visibility portion, leaving only need for compensation measures. Next, our sample sizes are too small to make these findings conclusive about all of OSS. In particular, we did not see differences between communities as we might have expected; breaking down our responses by artifacts and use cases did not show statistical differences between category means on any of the quantitative questions (0 of 16, 0%). Sampling OSS ecosystems broadly as we did here may prove less helpful going forward than conducting targeted studies to benefit particular communities, though. So future efforts might focus on specific artifacts, use cases, geographic regions, or programming language communities. Third, our sample of convenience via social media prevented us from engaging with those who use social media less frequently meaning our sample may contain biases toward social media users and more socially-engaged ecosystems.
Perhaps the greatest need arising from this work is a granular understanding of how specific artifacts facilitate the existence of labor data, its sharing, accessibility, and participant compensation. For example, we saw that certain attribution mediums corresponded to greater credit satisfaction. Further research might confirm whether appropriately selecting credit mediums improves accessibility of attribution and, in turn, improves worker satisfaction. These artifacts also shape worker perceptions of visibility, so researchers and practitioners likely need to pay careful attention to the interplay between accessibility and perceptions of openness. Labor will not become fully visible without ensuring consistent support for each factor that limits labor visibility and compensation, so articulation work and infrastructuring activities would benefit from renewed attention.
And of course, one study of invisibility will not make all labor visible, so future work might triangulate our findings within OSS through other qualitative and quantitative methods. Studies of invisible labor need not limit themselves to OSS contexts either as invisible labor happens everywhere, in computing and more broadly. For CSCW, this article provides further evidence in support of what Toxtli & colleagues (Toxtli et al., 2021) recently found in crowd work. We hope others will similarly build on our work in OSS and other domains.
6. Conclusion
Much of the work that happens in OSS ecosystems is not visible, not compensated, or both. Such invisible labor makes it difficult for individuals to advance their careers and for organizations to understand processes that enable their organizations to run. In this study, we measured invisible labor in OSS ecosystems by estimating the visibility and compensation rates of OSS activities. We found that approximately half of OSS labor does not receive credit, a low-cost form of attribution that often precedes other forms of compensation, and that OSS activities are more likely to be nonvisible or partially visible than to be fully visible. In addition, we demonstrated that anchoring labor to the concept of visibility tends to increase visibility frequency perceptions and decrease credit importance assessments. These findings tell us that compensating all those who work in OSS ecosystems will require careful attention to documenting labor activities, as well as sharing and making documentation accessible.
Acknowledgements.
Thanks to Jean-Gabriel Young and Melania Lavric for their suggestions throughout. The authors were partially supported by Google under the Open Source Complex Ecosystems and Networks project. J.M. was partially supported by the Sloan Foundation through the Vermont Research Open Source Program Office, and by the NSF award BCS-2332598.
Appendix A Invisible Labor Survey
The following are the complete survey questions as constructed in Qualtrics. A consent form (Q1) preceded these questions with the title “Open Source Credit Study” and an invitation to participate in a research study “about how people receive credit for the tasks they do as part of open source projects.” Italicized text indicates a question response constraint (e.g. a multiple choice question versus a multiple select question).
  1. Q1.
    Consent form (Click to consent)
  2. Q2.
    The questions in this survey ask about your experiences with open source projects. Specifically, the questions ask about projects you worked on during the last 2 years, and that at least one other person also worked on. (no response requested)
  3. Q3.
    The following categories describe things that projects either create or work on. Which categories apply to the projects you worked on? Select all that apply. (multiple select)
    • Hardware
    • Operating systems
    • Languages
    • Frameworks
    • Libraries
    • Packages
    • Applications
    • Data science
    • Scripts
    • Web-related
    • Something else (with text field)
  4. Q4.
    Projects create for different use cases. Which cases apply to the projects you worked on? Select all that apply. (multiple select)
    • Browsers
    • Business & productivity applications
    • Communication technologies
    • Content management
    • Education
    • Entertainment
    • Manufacturing
    • Publishing
    • Scientific research
    • Software infrastructure
    • Something else (with text field)
  5. Q5.
    Tasks vary in shape, size, and how visible they are to other people. For about 30 seconds, think through the different kinds of tasks you have done as part of these projects over the past 2 years. When you are ready, click on the right arrow to go on to the next question. (no response requested)
  6. Q6.
    The following questions ask about receiving credit. When we talk about having received credit for a task, we mean that the project told someone that you did something for the project. Those people may also work on the project or may be outside of the project. (no response requested)
  7. Q7.
    In the past two years, from how many projects did you receive credit for your tasks? (multiple choice)
    • 1 - None of them
    • 2 - Few of them
    • 3 - Some of them
    • 4 - Most of them
    • 5 - All of them
    • I’m not sure
  8. Q8.
    In the past two years, for how many of your tasks did you receive credit? (multiple choice)
    • 1 - None of them
    • 2 - Few of them
    • 3 - Some of them
    • 4 - Most of them
    • 5 - All of them
    • I’m not sure
  9. Q9.
    Projects give people credit for tasks through different mediums. Through what mediums did you receive credit? Select all that apply. (multiple select)
    • Automated metrics (e.g. GitHub Contributors)
    • Project membership (e.g. GitHub Organizations)
    • Project documentation (e.g. acknowledgement lists, release notes)
    • Standardized tools (e.g. All Contributors, CRediT)
    • Social media (e.g. Twitter, Facebook)
    • Blogs (e.g. Medium, personal websites)
    • Presentations (e.g. talks, conferences)
    • Something else (with text field)
    • I’m not sure
  10. Q10.
    How satisfied are you with the mediums through which you receive credit? (multiple choice)
    • 1 - Extremely dissatisfied
    • 2 - Dissatisfied
    • 3 - Neither satisfied nor dissatisfied
    • 4 - Satisfied
    • 5 - Extremely satisfied
  11. Q11.
    Think about all the tasks you performed again. (no response requested)
  12. Q12.
    (Randomized) How often did 2 or more people know that you performed those tasks? (multiple choice)
    • 1 - Never
    • 2 - Rarely
    • 3 - Sometimes
    • 4 - Often
    • 5 - Always
    • I’m not sure
  13. Q13.
    How often did 1 other person know that you performed those tasks? (multiple choice)
    • 1 - Never
    • 2 - Rarely
    • 3 - Sometimes
    • 4 - Often
    • 5 - Always
    • I’m not sure
  14. Q14.
    (Randomized) How often did nobody else know that you performed those tasks? (multiple choice)
    • 1 - Never
    • 2 - Rarely
    • 3 - Sometimes
    • 4 - Often
    • 5 - Always
    • I’m not sure
  15. Q15.
    How satisfied are you with how many of your tasks received credit? (multiple choice)
    • 1 - Extremely dissatisfied
    • 2 - Dissatisfied
    • 3 - Neither satisfied nor dissatisfied
    • 4 - Satisfied
    • 5 - Extremely satisfied
  16. Q16.
    How important is it to you to receive credit for the tasks you do? (multiple choice)
    • 1 - Not at all important
    • 2 - Slightly important
    • 3 - Moderately important
    • 4 - Very important
    • 5 - Extremely important
  17. Q17.
    Think about how the projects you worked on gave people credit. From your perspective, what worked well about how they gave people credit? (free response field)
  18. Q18.
    What did not work well about how they gave people credit? (free response field)
  19. Q19.
    Is there anything else that you’d like to share with us? (free response field)